OpenStudy (anonymous):

who could help me here out plz  8.Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.  9. A design on the surface of a balloon is 5 cm wide when the balloon holds 71 cm of air. How much air does the balloon hold when the design is 10 cm wide? Explain the method you use to find the amount of air.  

OpenStudy (jdoe0001):

@mikeiordan25 are you still there?

OpenStudy (anonymous):

yes im here @jdoe0001

OpenStudy (jdoe0001):

sorry I seem to be a bit lagged

OpenStudy (jdoe0001):

but anyhow, the 8) refers to a "right circular cylinder", that is the vase

OpenStudy (jdoe0001):

$$ \text{lateral suraface of a "right circular cylinder"}\\ 2\pi rh\\ \text{what you have is}\\ 3000=2\pi \color{red}{r}50\\ \text{solving for "r"}\\ \color{red}{r}=\cfrac{3000}{2\pi } $$ @mikeiordan25

OpenStudy (jdoe0001):

for 9) what you have is a "sphere volume", the sphere being the balloon

OpenStudy (anonymous):

thankx  6.  Jason designed an arch made of wrought iron for the top of a mall entrance. The 11 segments between the two concentric circles are each 1.25 m long. Find the total length of wrought iron used to make the structure. Round the answer to the nearest meter.    7.  Consider the prism shown below. a. Draw a net for the prism and label all dimensions. b. Use the net to find the surface area of the prism.  

OpenStudy (jdoe0001):

as far as 9) $$ \text{sphere volume formula is}\\ \cfrac{4}{3}\pi r^3\\ \text{what you have is }71^3 \text{ for the voume thus}\\ \text{you diameter is 5, half that is the radius}\\ 71^3=\cfrac{4}{3}\pi (2.5)^3\\ \text{now, what if the radius doubles?}\\ \text{that is, the diameter is 10, and radius 5?}\\ volume=\cfrac{4}{3}\pi (2\times 2.5)^3\\ $$

OpenStudy (jdoe0001):

just post the others in the channel, so we can all see it :)

OpenStudy (anonymous):

thankx @jdoe0001 and I have 6 and 7 my last plz

OpenStudy (jdoe0001):

\(71cm^3 \) I meant, since it's volume

OpenStudy (jdoe0001):

post 6 and 7 in the channel, so we can all see them :) notice you're missing a picture in 7) btw, don't forget to provide us that too

OpenStudy (anonymous):

ok @jdoe0001

OpenStudy (jdoe0001):

correction to above "r" formula :/ $$ \text{lateral surface of a "right circular cylinder"}\\ 2\pi rh\\ \text{what you have is}\\ 3000=2\pi \color{red}{r}50\\ \text{solving for "r"}\\ \color{red}{r}=\cfrac{3000}{2\pi 50 } $$